# LlamaRestTest: Effective REST API Testing with Small Language Models

### v1

@@ 4.7. Threats to Validity
A key concern in our study is the generalizability of the results, as we tested a limited range of RESTful services, including FDIC, Language-Tool, OhSome, OMDb, REST-Countries, Genome-Nexus, OCVN, Spotify, and YouTube-Mock API. These services may not represent all possible types of REST APIs, and different types of services could yield varying outcomes. Additionally, we relied on automated scripts for fault detection and code coverage, but potential bugs in these scripts or misalignment with API behavior could affect accuracy.
The metrics we used, such as fault detection, code coverage, and operation coverage, may not fully capture the effectiveness of the testing tools. High code coverage, for instance, doesn’t necessarily translate into better fault detection or practical bug identification. Our reliance on 500 status code errors as indicators of faults could also overlook more subtle bugs that don’t result in crashes. Despite repeating tests 10 times to mitigate randomness, variations in testing environments and API tools, especially in the computer setting, could still lead to inconsistencies.
@@ 5. Related Work
Automated REST API Testing: Automated testing for REST APIs has seen various strategies. White-box and black-box approaches, like those in EvoMaster, leverage evolutionary algorithms to refine test cases dynamically, with a focus on internal server errors (500 status code) detection (Arcuri, 2019). Black-box methodologies, exemplified by RESTler, generate stateful tests by deducing producer-consumer dependencies, aiming at detecting server failures (Atlidakis et al., 2019). RestTestGen and similar tools exploit data dependencies and employ oracles for validating status codes and response schemas (Corradini et al., 2022). MoRest (Liu et al., 2022) focuses on model-based testing to simulate user interactions and test case generation, while RestCT (Wu et al., 2022) employs combinatorial testing techniques to systematically explore parameter value combinations and their effects on API behavior. ARAT-RL introduces a reinforcement learning approach, dynamically refining API testing strategies by adapting to real-time feedback (Kim et al., 2023b). Techniques like QuickREST, Schemathesis, and RESTest use property-based testing and various oracles to ensure response compliance with OpenAPI or GraphQL specifications (Karlsson et al., 2020; Hatfield-Dodds and Dygalo, 2022; Martin-Lopez et al., 2021). Tools such as Dredd, fuzz-lightyear, and Tcases offer diverse testing capabilities, from comparing expected responses to detecting vulnerabilities and validating Swagger-based specifications (Apiary, 2023; Yelp, 2023; tcases, 2023).
Requirements-Based Test Case Generation: This domain includes tools like ucsCNL and methods utilizing UML Collaboration Diagrams to articulate use case specifications (Barros et al., 2011). Techniques like Requirements by Contracts and the SCENT-Method have proposed scenario-based approaches and custom languages to encapsulate functional requirements (Nebut et al., 2003; Ryser and Glinz, 1999a). SDL-based approaches and RTCM convert requirements into executable models or employ natural language processing for test generation (Tahat et al., 2001; Yue et al., 2015). These methods focus on the accurate reflection of intended requirements, while LlamaRestTest targets REST APIs with more formal specifications, facilitating a structured and adaptive fault-finding process during testing.
NLP in Test Generation: NLP has been instrumental in enhancing test generation from program specifications. Techniques vary from extracting OpenAPI parts from semi-structured documentation with machine learning to generating tests from natural language descriptions (Cao et al., 2017; Yang et al., 2018; Badri et al., 2004; Ryser and Glinz, 1999b). While effective in their own right, these methods differ from LlamaRestTest’s approach to extract rules, highlighting the distinct pathways in NLP application to API testing.
@@ 6. CONCLUSION AND FUTURE WORK
We introduced LlamaRestTest, a new REST API testing approach that leverages the capabilities of large language models through fine-tuning and quantization to improve testing effectiveness and efficiency. LlamaRestTest incorporates two key components, LlamaREST-IPD and LlamaREST-EX, which focus on inter-parameter dependency detection and realistic parameter value generation. Our evaluation, performed on 12 real-world REST services, shows that LlamaRestTest can outperform state-of-the-art tools such as RESTler, EvoMaster, MoRest, and ARAT-RL, achieving significantly higher code coverage and a larger number of processed operations. Additionally, our ablation study highlights the importance of each component, with LlamaREST-IPD and server response messages being critical to achieving the best results. Our results also show that fine-tuning significantly enhances LlamaRestTest’s performance, even outperforming larger models like RESTGPT. Quantization further improves the model’s efficiency without substantial loss in accuracy, making LlamaRestTest both powerful and practical for real-world use. In future work, we will explore ways to expand our approach’s capabilities to handle a wide variety of API types and refine its dynamic interaction with server responses for even more effective test input generation.
DATA AVAILABILITY
To facilitate the review process and support further research, all pertinent materials such as source code, datasets, and supplementary resources have been stored in an anonymized repository (Anonymous, 2023).

### v2

@@ 4.7. Threats to Validity
Like any empirical study, there are potential threats to the validity of our results. Our selection of REST APIs may not be representative and affect external validity. Although we used a diverse set of 12 real-world services, including services that have been used in prior evaluations of REST API testing tools, LlamaRestTest may perform differently on other benchmarks containing services with different characteristics. For instance, our reliance on specific fine-tuning datasets, such as Martin-Lopez’s IPD database and APIs-guru, may affect the performance of our fine-tuned models on other services.
There are also possible threats related to the experiment setup that could affect the accuracy of our results. First, we only compared with a limited number of tools. To address this threat, we considered the best-performing state-of-the-art REST API testing tools available at the time of this study. Second, the settings used for the testing tools and services might not be optimal. To mitigate this threat, we used the same settings used in a previous study (Kim et al., 2024a) and hosted each testing tool and service on separate machines to prevent interference. Furthermore, to account for randomness in tool results, we ran each tool 10 times on each service (this was done for all studies except the ablation study). Third, our use of automated scripts to measure fault detection and code coverage could introduce inaccuracies due to bugs in the scripts. However, we used scripts that have been used in prior studies (Kim, 2024) and double-checked them, in addition to manually spot-checking some of the results. Finally, LlamaRestTest’s integration with ARAT-RL creates tool dependency, in that potential issues in ARAT-RL could affect LlamaRestTest’s performance. Also in this case, we thoroughly tested and spot-checked implementation and results. We also made LlamaRestTest and the experiment dataset available in our artifact (Kim, 2023), which allows other researchers to further validate and extend our evaluation.
@@ 5. Related Work
Automated REST API Testing. Automated testing for REST APIs has seen various strategies. White-box and black-box approaches, such as those in EvoMaster, leverage evolutionary algorithms to refine test cases dynamically, with a focus on internal server errors (500 status code) detection (Arcuri, 2019). Black-box methodologies, exemplified by RESTler (Atlidakis et al., 2019), generate stateful tests by deducing producer-consumer dependencies, aiming at detecting server failures. RestTestGen (Corradini et al., 2022) and similar tools exploit data dependencies and employ oracles for validating status codes and response schemas. MoRest (Liu et al., 2022) focuses on model-based testing to simulate user interactions and test case generation, while RestCT (Wu et al., 2022) employs combinatorial testing techniques to systematically explore parameter value combinations and their effects on API behavior. ARAT-RL (Kim et al., 2024a) introduces a reinforcement learning approach, dynamically refining API testing strategies by adapting to real-time feedback. Techniques like QuickREST (Karlsson et al., 2020), Schemathesis (Hatfield-Dodds and Dygalo, 2022), and RESTest (Martin-Lopez et al., 2021) use property-based testing and various oracles to ensure response compliance with OpenAPI or GraphQL specifications. Tools such as Dredd (Apiary, 2023), fuzz-lightyear (Yelp, 2023), and Tcases (Kimbrough, 2023) offer diverse testing capabilities, from comparing expected responses to detecting vulnerabilities and validating Swagger-based specifications. More recently, two additional REST API testing techniques were proposed: DeepREST (Corradini et al., 2024) and AutoRestTest (Kim et al., 2025). DeepREST uses deep reinforcement learning to discover implicit API constraints, employing a single agent that learns operation orderings through a reward mechanism. AutoRestTest uses LLMs to generate parameter values while relying on the semantic property dependency graph for efficient dependency identification and multi-agent reinforcement learning for dynamic optimization across all testing steps. In contrast, LlamaRestTest employs fine-tuned LLMs for detecting IPDs and generating semantically valid input values, as well as leveraging reinforcement learning via its integration with ARAT-RL. It is worth noting that because DeepREST and AutoRestTest were developed concurrently with this work, we were unable to include them in our empirical comparisons.
LLM-based Test Generation. There is a growing body of work on leveraging LLMs for test generation (Wang et al., 2024), much of it focusing on unit testing. For example, AthenaTest (Tufano et al., 2020) uses a fine-tuned model for unit test generation, whereas TestPilot (Schäfer et al., 2023) synthesizes JavaScript and TypeScript tests. ASTER (Pan et al., 2025) implements an LLM-based pipeline for generating unit test for Java and Python. Unlike these techniques, LlamaRestTest focuses on REST API testing, employing two fined-tuned models for specific tasks together with analysis of dynamic server feedback.
LLM-based REST API Analysis. Recent work has also applied LLMs for REST API specification inference and API invocation. For instance, RESTSpecIT (Decrop et al., 2024) demonstrated capabilities for automatically inferring API specifications and conducting black-box testing with minimal user intervention. NESTFUL (Basu et al., 2024) created a benchmark for evaluating how LLMs handle complex nested API calls, highlighting current limitations in processing interconnected API operations. RestGPT (Song et al., 2023) introduced a coarse-to-fine online planning approach that enhances task breakdown and API selection processes. In contrast, LlamaRestTest leverages fine-tuned LLMs, server response messages, and reinforcement learning for improving REST API testing.
@@ 6. Conclusion and Future Work
We introduced LlamaRestTest, a new REST API testing approach that leverages the capabilities of large language models through fine-tuning and quantization to improve testing effectiveness and efficiency. LlamaRestTest incorporates two key components, LlamaREST-IPD and LlamaREST-EX, which focus on inter-parameter dependency detection and realistic parameter value generation. Our evaluation, performed on 12 real-world REST services, shows that LlamaRestTest can outperform state-of-the-art tools such as RESTler, EvoMaster, MoRest, and ARAT-RL, achieving significantly higher code coverage and processing more service operations than those tools. Additionally, our ablation study highlights the importance of each component, with LlamaREST-IPD and server response messages being critical to achieving the best results. Our results also show that fine-tuning significantly enhances LlamaRestTest’s performance, even outperforming larger models such as GPT. Quantization further improves the model’s efficiency without substantial loss in accuracy, making LlamaRestTest both powerful and practical for real-world use. In future work, we will explore ways to expand our approach’s capabilities to handle a wide variety of API types and refine its dynamic interaction with server responses for even more effective test input generation.
Acknowledgments
This work was partially supported by NSF, under grant CCF-0725202 and DOE, under contract DE-FOA-0002460, and gifts from Facebook, Google, IBM Research, and Microsoft Research.